OpenStudy (anonymous):

Solve the equation: (3t-5)/(t-1)=2+(2t)/(1-t)

myininaya (myininaya):

First we know t cannot =1 because if t=1 both sides are undefined. Also 1-t=-(t-1). So we have (3t-5)/(t-1)=2-2t/(t-1)

myininaya (myininaya):

to get rid of the t-1 in the denominator multiply both sides of the equation by t-1

myininaya (myininaya):

Now you have 3t-5=2(t-1)-2t

myininaya (myininaya):

Let me know if you need anymore help with this